solve the equation (x+6)(x-2)=(2x+1)(x-2)
I start this with creating trinomials from each side. x^2-2x+6x-12=2x^2-4x+1x-2. Now to get 0 on one side I Then subtract 2x^2, subtract 3x, add 2 from the right and place these values on the left. x^2-2x^2 +4x-3x-14 and combine like terms . I'm left with x^2+x-14=0. Now I have to break these two down as (x+ )(x- ) so I have a solution set. hELP I'm STUMPED what can I do that would leave me with 1x in the middle when I only can use 4X4 or 7X2?

Answer by richwmiller(17219)


x^2+4x-12 = 2x^2-3 x-2
-x^2+7 x-10 = 0
-(x-5) (x-2) = 0
x=2
x=5